OpenStudy (anonymous):

2(3y-9) at least or equal to (2y +5)

sam (.sam.):

2(3y-9)<=(2y+5) 2(3y-9)<=2y+5 6y-18<=2y+5 4y-18<=5 4y<=23 y<=(23)/(4)